### v1

# General Perturbative Expansion and Relations with Unitarity Triangles

Abstract
In this letter, using a rephasing invariant formula $\delta=\arg[{V_{ud}V_{us}V_{cb}V_{tb}/V_{ub}\det V_{\rm CKM}}]$, we evaluate the CP phase $\delta$ of the CKM matrix $V_{\rm CKM}$ perturbatively for small quark mixing angles $s_{ij}^{u,d}$ with associated phases $\rho_{ij}^{u,d}$. Consequently, we derived a relation $\delta\simeq\arg[\Delta s_{12}\Delta s_{23}/(\Delta s_{13}-s^{u}_{12}e^{-i\rho^{u}_{12}}\Delta s_{23})]$ with $\Delta s_{ij}\equiv s^{d}_{ij}e^{-i\rho^{d}_{ij}}-s^{u}_{ij}e^{-i\rho^{u}_{ij}}$. Such a result represents the analytic behavior of the CKM phase. The uncertainty in the relation is of order $O(\lambda^{2})\sim 4\%$, which is comparable to the current experimental precision. Comparisons with experimental data suggest that the hypothesis of some CP phases being maximal. We also discussed relationships between the phase $\delta$ and unitarity triangles. As a result, several relations between the angles $\alpha,\beta,\gamma$ and $\delta$ are identified through other invariants $V_{il}V_{jm}V_{kn}/\det V_{\rm CKM}$.
preprint: STUPP-25-283
I Introduction
Understanding the origin of CP violation is crucial for explaining the baryon asymmetry in the universe. CP violation in the CKM matrix Kobayashi:1973fv has been discussed in various forms throughout the history of particle physics Fritzsch:1995nx ; Rasin:1997pn ; Harrison:2009bz ; Xing:2020ijf . In particular, Ref. Hall:1993ni presented a general treatment of the CKM matrix and its CP-violating phase. However, such general treatments have not been pursued in later literature. Furthermore, most analyses often use the Jarlskog invariant Jarlskog:1985ht , which is of order $O(10^{-5})$ in the quark sector. Since the small invariant is highly sensitive to various approximations, capturing its analytic behavior—including error estimation—has been a technically challenging task from the standpoint of computational precision.
In this letter, employing a recently proposed rephasing invariant formula Yang:2025hex , we derive a general perturbative expression for the CP-violating phase in the CKM matrix, and its phenomenological consequences.
II A Rephasing Invariant formula for CP Phase of CKM Matrix
and Its Perturbative Expansion
We begin by presenting a method to directly extract the CP phase $\delta$ from the CKM matrix $V_{\rm CKM}$ defined in a general phase basis. To convert a given $V_{\rm CKM}$ into the PDG standard parametrization $V_{\rm CKM}^{0}$, we remove unphysical phases by applying redefinition of phases as
$V_{\rm CKM}^{0}=\Psi_{L}^{\dagger}V_{\rm CKM}\Psi_{R}\,.$ (1)
Here, $\Psi_{L,R}={\rm diag}(e^{i\gamma_{(L,R)1}},\,e^{i\gamma_{(L,R)2}},\,e^{i\gamma_{(L,R)3}})$ are diagonal phase matrices, and each $\gamma_{(L,R)i}$ represents an arbitrary phase. Due to the overall phase redundancy, the number of independent degrees of freedom is five.
In the standard PDG parametrization, elements of the mixing matrix satisfy the following conditions:
$\arg V_{ud}^{0}=\arg V_{us}^{0}=\arg V_{cb}^{0}=\arg V_{tb}^{0}=0\,,~~\arg\left[V_{cd}^{0}V_{ts}^{0}-V_{cs}^{0}V_{td}^{0}\right]=\arg\left[V_{ub}^{0*}\det V_{\rm CKM}^{0}\right]=\delta\,.$ (2)
The final constraint follows from the identity for the inverse of a unitary matrix.
Using these five conditions, one can solve for the five phase parameters $\gamma_{(L,R)i}$. As a result, for a CKM matrix given in an arbitrary basis of phases, the CP-violating phase $\delta$ in the PDG convention is expressed as
$\delta$ $=\arg\left[\frac{V_{ud}V_{us}V_{cb}V_{tb}}{V_{ub}\det V_{\rm CKM}}\right]\,.$ (3)
This expression is explicitly rephasing invariant, including the phase of $\det V_{\rm CKM}$ Wu:1985ea ; Bjorken:1987tr ; Nieves:1987pp ; Jenkins:2007ip , and it clearly coincides with the phase $\delta$ in the PDG standard parametrization.
The formula possesses several advantages over the traditional Jarlskog invariant Jarlskog:1985ht , as summarized below:
1. Factorizability:
It can be decomposed into individual elements $V_{\alpha\beta}$ and the determinant $\det V_{\rm CKM}$, making the computation more transparent and straightforward.
2. Robustness under approximations:
The Jarlskog invariant $J$ is highly sensitive to approximations because of its smallness $O(10^{-5})$ in the quark sector. On the other hand, the new invariant is of order $O(1)$ and therefore less affected by perturbative corrections. Furthermore, its uncertainty is easier to quantify.
3. Completeness of phase information:
The invariant $J$ does not retain the sign of $\cos\delta$, requiring additional calculations to reconstruct the full experimental results. In contrast, this formula directly preserves the full information on the CP-violating phase $\delta$.
We now demonstrate the relation between this formula and the well-known Jarlskog invariant. By dividing the complex quantity (inside the argument) by its modulus, the phase is explicitly extracted as
$e^{i\delta}=\frac{V_{ud}V_{us}V_{cb}V_{tb}}{V_{ub}\det V_{\rm CKM}}\left|\frac{V_{ub}\det V_{\rm CKM}}{V_{ud}V_{us}V_{cb}V_{tb}}\right|=\frac{V_{ud}V_{us}V_{cb}V_{tb}V_{ub}^{*}\det V_{\rm CKM}^{*}}{|V_{ud}V_{us}V_{cb}V_{tb}V_{ub}|}\,.$ (4)
Here, we used the identity $\det V_{\rm CKM}\det V_{\rm CKM}^{*}=|\det V_{\rm CKM}|^{2}=1$. Since our goal is to obtain the invariant $V_{ud}V_{tb}V_{ub}^{*}V_{td}^{*}$, An alternative element of the inverse matrix $V_{us}\det V_{\rm CKM}^{*}=V_{cb}^{*}V_{td}^{*}-V_{cd}^{*}V_{tb}^{*}$ yields
$e^{i\delta}=\frac{V_{ud}V_{cb}V_{tb}V_{ub}^{*}(V_{cb}^{*}V_{td}^{*}-V_{cd}^{*}V_{tb}^{*})}{|V_{ud}V_{us}V_{cb}V_{tb}V_{ub}|}=\frac{|V_{cb}|^{2}V_{ud}V_{tb}V_{ub}^{*}V_{td}^{*}-|V_{tb}|^{2}V_{ud}V_{cb}V_{ub}^{*}V_{cd}^{*}}{|V_{ud}V_{us}V_{cb}V_{tb}V_{ub}|}\,.$ (5)
Taking the imaginary parts of both sides, the right-hand side contains the Jarlskog invariant $J\equiv{\rm Im}\,[V_{ud}V_{tb}V_{ub}^{*}V_{td}^{*}]$. Using the orthonormal relation $|V_{cb}|^{2}+|V_{tb}|^{2}=1-|V_{ub}|^{2}$, we obtain
${\rm Im}\,e^{i\delta}=\frac{(1-|V_{ub}^{2}|)J}{|V_{ud}V_{us}V_{cb}V_{tb}V_{ub}|}=\sin\delta\,.$ (6)
The last equality follows directly from an expression $J=c_{12}s_{12}c_{23}s_{23}c_{13}^{2}s_{13}\sin\delta$ with the observed mixing angles $s_{ij},c_{ij}$. Therefore, this $\sin\delta$ coincides with the value derived from the Jarlskog invariant.
Next, we perform the perturbative expansion. The CKM matrix $V_{\rm CKM}\equiv U_{u}^{\dagger}U_{d}$ is defined as the misalignment between the diagonalization matrices of the left-handed up-type quarks $U_{u}$ and down-type quarks $U_{d}$. By choosing an appropriate basis, elements of both $U_{u,d}$ are taken to be of the same order as those of $V_{\rm CKM}$ without loss of generality. Therefore, we adopt the following approximation.
Approximation:
The mixing angles $s_{ij}^{u,d}\equiv\sin\theta_{ij}^{u,d}\,,\,c_{ij}^{u,d}\equiv\cos\theta_{ij}^{u,d}$ of $U_{u,d}$ are assumed to satisfy $s_{12}^{u,d}\sim\lambda$, $s_{23}^{u,d}\sim\lambda^{2}$, and $s_{13}^{u,d}\sim\lambda^{3}$, with an expansion parameter $\lambda\simeq 0.2$.
Justification:
When the Yukawa matrices $Y_{u,d}$ of quarks possess chiral symmetries for the first and second generations, $Y_{u,d}=D_{L}Y_{u,d}D_{R}$, all lighter singular values and mixings vanish. Here, $D_{L,R}\equiv{\rm diag}(e^{i\phi_{L,R}^{1}},e^{i\phi_{L,R}^{2}},1)$ and $\phi_{L,R}^{1,2}$ are phases. Although these chiral symmetries are only approximate in reality, the mixing angles are suppressed by powers of corresponding ratios of singular values $m_{fi}/m_{fj}$.
We now proceed to define the notation of perturbative expansion. The matrices $U_{u,d}$ are generally written as $U_{u,d}=\Phi_{u,d}^{L}U_{u,d}^{0}\Phi_{u,d}^{R}$ with diagonal phase matrices $\Phi_{u,d}^{L,R}$ and their PDG parametrizations $U_{u,d}^{0}$. Due to the freedom of right-handed phase transformations, the unitary matrices are redefined as $U_{u,d}^{1}=\Phi_{u,d}^{L}U_{u,d}^{0}\Phi_{u,d}^{L\dagger}$ without loss of generality. Since $U_{u,d}^{0}$ contain only small mixing angles, the leading order of the perturbation is approximated as follows:
$U_{u,d}^{1}\simeq\begin{pmatrix}1&s_{12}^{u,d}e^{-i\rho_{12}^{u,d}}&s_{13}^{u,d}e^{-i\rho_{13}^{u,d}}\\ -s_{12}^{u,d}e^{i\rho_{12}^{u,d}}&1&s_{23}^{u,d}e^{-i\rho_{23}^{u,d}}\\ -s_{13}^{u,d}e^{i\rho_{13}^{u,d}}+s_{12}^{u,d}s_{23}^{u,d}e^{i\rho_{12}^{u,d}+i\rho_{23}^{u,d}}&-s_{23}^{u,d}e^{i\rho_{23}^{u,d}}&1\\ \end{pmatrix},$ (7)
where $\rho_{ij}^{u,d}$ are the associated CP-violating phases corresponding to the mixing angles. The next-to-leading order terms in each matrix element are suppressed by at least order $\lambda^{2}$ compared to the leading order. Since the right-handed phases $\Phi^{R}_{u,d}$ of quarks do not affect the observed CP phase, we will omit them hereafter.
In this case, the CKM matrix to be analyzed is redefined as $V_{\rm CKM}\equiv U_{u}^{1\dagger}U_{d}^{1}$. Expanding arguments of each matrix element in powers of $\lambda$, we obtain
$\arg V_{ud}$ $=0+O(\lambda^{2})\,,~~~\arg V_{tb}=0+O(\lambda^{4})\,,$ (8)
$\arg V_{us}$ $=\arg\left[s^{d}_{12}e^{-i\rho^{d}_{12}}-s^{u}_{12}e^{-i\rho^{u}_{12}}\right]+O(\lambda^{2})\,,~~~\arg V_{cb}=\arg\left[s^{d}_{23}e^{-i\rho^{d}_{23}}-s^{u}_{23}e^{-i\rho^{u}_{23}}\right]+O(\lambda^{2})\,,$ (9)
$\arg V_{ub}$ $=\arg\left[s^{d}_{13}e^{-i\rho^{d}_{13}}-s^{u}_{13}e^{-i\rho^{u}_{13}}-s^{u}_{12}e^{-i\rho^{u}_{12}}(s^{d}_{23}e^{-i\rho^{d}_{23}}-s^{u}_{23}e^{-i\rho^{u}_{23}})\right]+O(\lambda^{2})\,.$ (10)
In the limit $s_{ij}^{u,d}\to 0$, the associated CP phases $\rho_{ij}^{u,d}$ simultaneously vanish. Therefore, the contributions from these CP phases appear at first order in the mixing angles. Note that observed mixing angles $s_{ij}$ and $c_{ij}$ constraint the absolute values of matrix elements as,
$|V_{us}|=s_{12}c_{13}\,,~~|V_{ub}|=s_{13}\,,~~|V_{cb}|=s_{23}c_{13}\,.$ (11)
From the rephasing invariant formula, a general perturbative relation for the CP phase of the CKM matrix will be
$\delta$ $=\arg[V_{us}V_{cb}/V_{ub}]+O(\lambda^{2})\,,$ (12)
$\left|V_{us}\over V_{ub}/V_{cb}\right|e^{i\delta}$ $=\frac{s^{d}_{12}e^{-i\rho^{d}_{12}}-s^{u}_{12}e^{-i\rho^{u}_{12}}}{\dfrac{s^{d}_{13}e^{-i\rho^{d}_{13}}-s^{u}_{13}e^{-i\rho^{u}_{13}}}{s^{d}_{23}e^{-i\rho^{d}_{23}}-s^{u}_{23}e^{-i\rho^{u}_{23}}}-s^{u}_{12}e^{-i\rho^{u}_{12}}}+O(\lambda^{2})\,.$ (13)
Since errors of this expression are $O(\lambda^{2})\simeq 4\%$, which is comparable to the current experimental uncertainties, the expression is sufficiently accurate. In particular, the absolute value of the denominator has a fixed value,
$\left|\dfrac{s^{d}_{13}e^{-i\rho^{d}_{13}}-s^{u}_{13}e^{-i\rho^{u}_{13}}}{s^{d}_{23}e^{-i\rho^{d}_{23}}-s^{u}_{23}e^{-i\rho^{u}_{23}}}-s^{u}_{12}e^{-i\rho^{u}_{12}}\right|=\left|V_{ub}\over V_{cb}\right|=0.09\,.$ (14)
Thus, depending on the magnitude of $s_{12}^{u}$, one of the terms in the denominator can be neglected.
It is theoretically intriguing to investigate origins of the observed large CP phase. In situations where the 1-3 mixing angles $s_{13}^{u,d}$ are sufficiently smaller than $s_{12}^{u}V_{cb}$ by the above chiral symmetries, neglect of these terms yields
$\left|V_{us}\over V_{ub}/V_{cb}\right|e^{i\delta}$ $\simeq\frac{s^{d}_{12}e^{-i\rho^{d}_{12}}-s^{u}_{12}e^{-i\rho^{u}_{12}}}{-s^{u}_{12}e^{-i\rho^{u}_{12}}}+O(\lambda^{2})\,,$ (15)
$\delta$ $\simeq\arg\left[1-{s^{d}_{12}\over s^{u}_{12}}e^{i(\rho^{u}_{12}-\rho^{d}_{12})}\right]+O(\lambda^{2})\,.$ (16)
The observed physical value of CKM matrix in the latest UTfit is UTfit:2022hsi
$\sin\theta_{12}^{\rm CKM}$ $=0.22519\pm 0.00083\,,~~~\sin\theta_{23}^{\rm CKM}=0.04200\pm 0.00047\,,$ (17)
$\sin\theta_{13}^{\rm CKM}$ $=0.003714\pm 0.000092\,,~~~\delta_{q}=1.137\pm 0.022\,.$ (18)
Given this experimental value, the hypothesis that the phase is maximal $\rho^{d}_{12}-\rho^{u}_{12}=\pi/2$ appears quite plausible Shin:1985cg ; Gronau:1985tx ; Fritzsch:1985yv ; Kang:1985nw ; Lehmann:1995br ; Kang:1997uv ; Fritzsch:1999im ; Antusch:2009hq ; Xing:2003yj ; Yang:2020qsa ; Yang:2020goc ; Yang:2021smh . A non-perturbative treatments beyond perturbation theory for $s_{12}^{f}$ and $s_{23}^{f}$ are also available in previous papers Yang:2024ulq ; Yang:2025yst .
Relation between the phase and unitarity triangles
It is of particular interest to discuss relations between the results and unitarity triangles Harrison:2006bj ; Frampton:2010ii ; Dueck:2010fa ; Li:2010ae . From the alternative $s$-$b$ unitarity triangle, given by $V_{us}V_{ub}^{*}+V_{cs}V_{cb}^{*}+V_{ts}V_{tb}^{*}=0$, its three angles are rephasing invariant quantities,
$\alpha^{\prime}=\arg\left[-{V_{ts}V_{tb}^{*}\over V_{us}V_{ub}^{*}}\right],~~\beta^{\prime}=\arg\left[-{V_{cs}V_{cb}^{*}\over V_{ts}V_{tb}^{*}}\right],~~\gamma^{\prime}=\arg\left[-{V_{us}V_{ub}^{*}\over V_{cs}V_{cb}^{*}}\right].$ (19)
If the phase of $V_{cs}$ can be neglected, the phase $\delta$ is related to one of angles of the unitarity triangle.
$\gamma^{\prime}+\pi=\arg\left[{V_{us}V_{cb}\over V_{ub}V_{cs}}\right]\simeq\delta\,.$ (20)
Indeed, relations between the phase and the angles are
$\delta+\alpha^{\prime}=\arg\left[\frac{V_{ud}V_{us}V_{cb}V_{tb}}{V_{ub}\det V_{\rm CKM}}\right]+\arg\left[-{V_{ts}V_{tb}^{*}\over V_{us}V_{ub}^{*}}\right]=\arg\left[-\frac{V_{ud}V_{cb}V_{ts}}{\det V_{\rm CKM}}\right]=1.05^{\circ}\,,$ (21)
$\delta-\gamma^{\prime}-\pi=\arg\left[\frac{V_{ud}V_{us}V_{cb}V_{tb}}{V_{ub}\det V_{\rm CKM}}\right]-\arg\left[{V_{cb}V_{us}\over V_{ub}V_{cs}}\right]=\arg\left[\frac{V_{ud}V_{cs}V_{tb}}{\det V_{\rm CKM}}\right]=-0.0019^{\circ}\,,$ (22)
which define other rephasing invariants. Moreover, the angles of the standard unitarity triangle are
$\alpha=\arg\left[-{V_{td}V_{tb}^{*}\over V_{ud}V_{ub}^{*}}\right]=92.40^{\circ},~~\beta=\arg\left[-{V_{cd}V_{cb}^{*}\over V_{td}V_{tb}^{*}}\right]=22.49^{\circ},~~\gamma=\arg\left[-{V_{ud}V_{ub}^{*}\over V_{cd}V_{cb}^{*}}\right]=65.11^{\circ},$ (23)
and exact relations with these angles are found to be
$\delta+\alpha+\pi$ $=\arg\left[\frac{V_{ud}V_{us}V_{cb}V_{tb}}{V_{ub}\det V_{\rm CKM}}\right]+\arg\left[{V_{td}V_{tb}^{*}\over V_{ud}V_{ub}^{*}}\right]=\arg\left[\frac{V_{us}V_{cb}V_{td}}{\det V_{\rm CKM}}\right]=-22.45^{\circ}\fallingdotseq-\beta\,,$ (24)
$\delta-\gamma-\pi$ $=\arg\left[\frac{V_{ud}V_{us}V_{cb}V_{tb}}{V_{ub}\det V_{\rm CKM}}\right]+\arg\left[{V_{ud}V_{ub}^{*}\over V_{cd}V_{cb}^{*}}\right]=\arg\left[\frac{V_{us}V_{cd}V_{tb}}{\det V_{\rm CKM}}\right]=-179.96\fallingdotseq-\pi.$ (25)
These two expressions are not independent; from $\arg 2\pi=0$, we obtain
$\epsilon\equiv\delta+\alpha+\beta+\pi=\delta-\gamma=\arg\left[-\frac{V_{us}V_{cd}V_{tb}}{\det V_{\rm CKM}}\right]=0.035^{\circ}\,.$ (26)
Since the perturbative expansions of $V_{us}$ and $V_{cd}$ agree up to $\mathcal{O}(\lambda^{3})$ in Eq. (9), the phase of $-V_{us}V_{cd}V_{tb}$ is of order $\mathcal{O}(\lambda^{4})\sim 0.2\%$. At this point, it seems meaningful to examine the remaining two invariants;
$\arg\left[\frac{V_{ub}V_{cs}V_{td}}{\det V_{\rm CKM}}\right]=-87.60^{\circ}\fallingdotseq\alpha-\pi\,,~~~\arg\left[\frac{V_{us}V_{cd}V_{tb}}{\det V_{\rm CKM}}\right]=-64.01^{\circ}\simeq-\gamma\,.$ (27)
That is, three of the six invariants are approximately identified with the angles $\alpha-\pi$, $-\beta$, and $-\gamma$, and the remaining three represent small differences between the angles and invariants. Therefore, they provide an alternative perspective on the characteristic CP phases in the CKM matrix. A similar argument can be applied to other unitarity triangles.
III Summary
In this letter, using a rephasing invariant formula $\delta=\arg[{V_{ud}V_{us}V_{cb}V_{tb}/V_{ub}\det V_{\rm CKM}}]$, we evaluate the CP phase $\delta$ of the CKM matrix $V_{\rm CKM}$ perturbatively for small quark mixing angles $s_{ij}^{u,d}$ with associated phases $\rho_{ij}^{u,d}$. Consequently, we derived a relation $\delta\simeq\arg[\Delta s_{12}\Delta s_{23}/(\Delta s_{13}-s^{u}_{12}e^{-i\rho^{u}_{12}}\Delta s_{23})]$ with $\Delta s_{ij}\equiv s^{d}_{ij}e^{-i\rho^{d}_{ij}}-s^{u}_{ij}e^{-i\rho^{u}_{ij}}$. Such a result represents the analytic behavior of the CKM phase. The uncertainty in the relation is of order $O(\lambda^{2})\sim 4\%$, which is comparable to the current experimental precision. Comparisons with experimental data suggest that the hypothesis of some CP phases being maximal.
We also discussed relationships between the phase $\delta$ and unitarity triangles. As a result, several relations between the angles $\alpha,\beta,\gamma$ and $\delta$ are identified through other invariants $V_{il}V_{jm}V_{kn}/\det V_{\rm CKM}$. These general perturbative relations broadly cover phenomenological calculations, and therefore, the presented results have wide applicability in studies of flavor physics and CP violation.
Acknowledgment
The study is partly supported by the MEXT Leading Initiative for Excellent Young Researchers Grant Number JP2023L0013.